# DPZV: Elevating the Tradeoff between Privacy and Utility in Zeroth-Order Vertical Federated Learning

### v2

@@ 5 Convergence Analysis
Discussion. The first term is influenced by the model’s initialization, $F^{0}$. This term also enjoys the same rate as ZO optimization in the centralized case [14]. The second term $\mathcal{O}(\frac{d_{*}}{T})$ is a standard term for ZOO methods based on the usage of ZO estimator updates. The third term captures the impact of various noise sources in the learning system. Here, $\sigma_{s}^{2}/B$ represents the noise introduced by stochastic gradients, and $\sigma_{dp}^{2}$ corresponds to the variance of the injected DP noise. While reducing $\sigma_{dp}^{2}$ improves utility, it comes at the cost of weakening privacy guarantees. Consequently, this term encapsulates the fundamental trade-off between model performance, computational cost, and privacy budget. The fourth term, quantifies the impact of the gradient clipping operation on convergence. As $C$ increases, increases, this term diminishes, effectively recovering the non-private case in the limit. However, the noise variance $\sigma_{dp}^{2}$ also scales linearly with $C$ , which can impede overall convergence. This trade-off underscores the importance of carefully tuning the sensitivity level to balance privacy preservation and learning efficiency.
Based on the privacy analysis in Section 6, we further substitute the level of DP variance $\sigma_{dp}=\frac{2C\sqrt{T}}{D\mu}$ into this convergence rate. By selecting $T={\mathcal{O}}\left(\frac{D^{2}\mu^{2}\sqrt{d}}{C^{2}}\left({\mathcal{F}}+\frac{\sigma^{2}}{B}+\frac{C^{2}}{e^{C^{2}}-1}\right)\right),$ the overall convergence upper bound becomes ${\mathcal{O}}\left(\frac{\sqrt{d}}{D\mu}\right),$ which matches the known rate of DP-SGD for first-order methods [5]. This result underscores a key insight of our analysis: although zeroth-order optimization typically converges more slowly than its first-order counterpart in non-private settings, it can achieve the same convergence rate when subject to differential privacy constraints.
6 Privacy Analysis
While data are kept local in the VFL framework, the communication of local embeddings and backward gradient during the training process, as well as the sharing of model weights post training, can pose threat to sensitive information such as label and features [29]. In this section, we talk about how our algorithm protect privacy under these threat models. We consider two scenarios: 1) Honest-but-curious: In the honest-but-curious threat model, participants adhere strictly to the protocols of VFL without deviating from agreed procedures. However, they may attempt to infer private information from intermediate results exchanged during training. 2) Post-training adversarial: This model focuses on attacks conducted after the training process. It assumes the attacker lacks access to intermediate results. The adversary only has access to the trained model and aims to infer sensitive information by analyzing the model’s behavior or responses.
Theorem 6.1.
Under 5.1-5.3, suppose the privacy parameter $\sigma_{dp}$ is $\sigma_{dp}=\frac{2C\sqrt{T}}{D\mu},$ where $D$ denotes the volume of the dataset, $T$ defines total iterations, and $\mu>0$ controls the privacy level. The training process of Algorithm  1 is seen to be $(\epsilon,\delta(\epsilon))$-differential private for $\forall\epsilon>0$, where
$\textstyle\delta(\epsilon)=\Phi(-\frac{\epsilon}{\mu}+\frac{\mu}{2})-e^{\epsilon}\Phi(-\frac{\epsilon}{\mu}-\frac{\mu}{2})$ (11)
Discussion. The theorem provides privacy guarantee under the "honest-but-curious" threat model, where one or a few malicious clients try to do inference attacks by collecting information of the system. By providing only the differentially private ZO information $\Delta_{m}^{t}$, the algorithm protects for both label inference attacks [13] and feature inference attacks [25], because the attacker cannot differentiate a single datum in the dataset ${\mathcal{D}}$ and label set ${\bm{Y}}$. The differential privacy parameter $\mu$ is related to $(\epsilon,\delta)$ through the relationship defined in (11). Given any two of these parameters, the third can be determined by solving (11), allowing full flexibility in controlling the privacy level.
Although the forward embeddings $h_{m,i}$ are not protected by differential privacy and may be vulnerable to feature attacks from the server, existing attacks primarily rely on access to model parameters $w_{m}$ for model inversion [16, 18] or require backward gradients for gradient inversion [19]. Both attack vectors are mitigated under our differential privacy guarantees. A detailed discussion of our framework’s robustness against various attacks is provided in Appendix E, and the detailed proofs can be found in Appendix C.
7 Experiments
Setup. We consider four datasets: image dataset MNIST [7], CIFAR-10 [20], semantic dataset Amazon Review Polarity [27], and multi-view dataset ModelNet40 [34]. For each dataset, we conduct a grid search on learning rates and other hyperparameters. We run 100 epochs on each method and select the best validation model. We run each algorithm under three random seeds and compute the sample variance for the generosity of our result. For MNIST, we use a two-layer CNN model, for VFL’s data partitioning, we split the images by row evenly into 7 sub-images and assign them to 7 clients. For CIFAR-10 we use a four-layer CNN model, and partition each image into 2×2, 4 patches of the same size for 4 clients. For ModelNet40, we use a ResNet-18 model, and partition each object into 12 different camera views and allocate them to 12 clients. For Amazon Reviews, we use a pre-trained BERT [8] model, and split the tokenized data input into 3 paragraphs of the same number of tokens and distributed them across 3 clients. For all four datasets, we use a fully connected model of two linear layers with ReLU activations as the server model.
We compare our algorithm against several SotA VFL methods: 1) VAFL [4] 2) ZOO-VFL [39], 3) VFL-CZOFO [33]. All methods assume that the server holds the labels, and concatenates the embeddings of clients as the input of the server. VAFL updates its model through first-order optimization in an asynchronous manner, and achieves DP by adding random noise to the output of each local embedding. We use VAFL as a first-order baseline of VFL method. Contrary to VAFL, ZOO-VFL and ZOFO both adopt ZO optimization in their training procedure. ZOO-VFL substitutes the first order optimization by ZO optimization in common VFL methods. VFL-CZOFO uses a cascade hybrid optimization method that computes the intermediate gradient via ZOO, while keeping the back propagation on both server and client. To enforce DP and compare all methods fairly, we follow the approach in VAFL [4], applying the same DP mechanism to both VFL-CZOFO and ZOO-VFL, which involves clipping the embeddings and adding calibrated vector noise. While VAFL updates server and client once for every communication round, some first-order VFL methods adopt multiple local updates in one communication round [24, 3]. In this paper, we only focus on the same update behavior as VAFL. In addition, model delay has been manually adjusted for all methods based on the per-batch computation time on a single client to simulate client heterogeneity and ensure a fair comparison.
Comparison with Baselines. Figure 2 presents the performance evaluation of DPZV against all baselines on image classification tasks under certain DP levels. The results show that DPZV consistently outperforms all baselines on MNIST and CIFAR-10 under strict privacy budget ($\epsilon=1,\delta=1\times 10^{-3}$). On ModelNet40, where training is more challenging due to larger models and a greater number of clients, we slightly relax the DP constraint ($\epsilon=10$), and observe that DPZV maintains a competitive advantage over other ZO-based methods while achieving accuracy comparable to the first-order baseline VAFL. The scalar noise injected in DPZV effectively constrains the noise magnitude, mitigating the instability typically introduced by noisy zeroth-order gradient estimators. This leads to more stable training dynamics than other ZO based methods. Moreover, in contrast to the high-dimensional vector noise used in first-order baseline VAFL, scalar noise incurs significantly less loss, resulting in superior performance under stringent privacy budgets.
Privacy-Accuracy Tradeoff. Figure 3 presents the accuracy-privacy tradeoff across four benchmark datasets: MNIST, CIFAR-10, ModelNet40, and Amazon Reviews. We evaluate the robustness of our proposed method DPZV under various privacy budgets $\epsilon$, while fixing the failure probability $\delta=1\times 10^{-3}$. Across all tasks, DPZV consistently achieves the highest accuracy under tight privacy regimes ($\epsilon\leq 1$), indicating its ability to maintain model utility even with stringent differential privacy constraints. Notably, on MNIST and ModelNet40, DPZV shows only a marginal drop in accuracy as $\epsilon$ decreases, while the competing methods suffer substantial degradation. For instance, on CIFAR-10 at $\epsilon=0.1$, DPZV maintains over 90% accuracy, whereas ZOFO and ZOO-VFL fall below 30%. Similar trends are observed on Amazon Reviews, where DPZV consistently outperforms baselines under strict privacy, especially at $\epsilon=1$ and $\epsilon=5$. These results validate the effectiveness of our scalar-noise-based DP mechanism in balancing utility and privacy, and highlight its advantage in real-world privacy-sensitive federated learning scenarios.
Effect of Clipping in Non-Private Scenarios. We observe in Figure 3 that even under no privacy constraints ($\epsilon=\infty$), DPZV outperforms other ZO based algorithms. A key distinction lies in DPZV’s use of scalar clipping on ZO information, while originally introduced for differential privacy, also acts as a form of gradient regularization. This regularization effect has been shown to improve convergence in prior work [37], and we observe similar benefits here. Figure 5 verifies our insight by comparing the convergence behavior of DPZV with and without gradient clipping under a non-private setting. The plots show test loss versus communication rounds. In both datasets, applying clipping to the ZO information significantly improves convergence speed. For MNIST, clipped DPZV reaches low test loss much faster and stabilizes more smoothly. On CIFAR-10, the clipped version also demonstrates consistently lower loss throughout training. These results suggest that clipping not only stabilizes training but also enhances convergence efficiency, even when privacy is not enforced.
8 Conclusion and Limitations
In this work, we propose DPZV, the first zero-order VFL framework that achieves robust and controllable privacy guarantees while maintaining strong model utility. Through rigorous theoretical analysis, we establish the strong convergence properties of our algorithm. Extensive experiments demonstrate that DPZV outperforms all baselines across various privacy levels and cost evaluations. Our method also features a lightweight memory footprint, enabling enhanced performance on large language models. These results provide valuable insights for future advancements in VFL and distributed ZO optimization methods.
We identify several potential limitations of our work. Notably, the convergence rate established in our theoretical analysis remains dimension-dependent, consistent with prior results in centralized ZO optimization. Extending the dimension-free convergence guarantees shown in [26] to the VFL setting poses a nontrivial challenge due to the inherently distributed and feature-partitioned nature of VFL. Investigating such extensions is a promising direction for future work.
Appendix
[…]
Appendix A Preliminaries and Outline
We first define the following notation table to facilitate the proof:
Notation Description
${\bm{w}}=[w_{0},w_{1},w_{2},\dots,w_{M}]$ All learnable parameters
$d_{0},d_{1},\dots,d_{M}$ Dimension of parameters on server (machine 00) and client $1,\dots,M$
$d=\sum_{m=0}^{M}d_{m}$ Dimension of all parameters
$F({\bm{w}}):=F({\bm{w}};{\mathcal{D}},{\bm{Y}})$ Global loss function
$\chi^{t}=[w_{1}^{t_{1}},\dots,w_{M}^{t_{M}}]$ Latest learnable parameters of all clients at server time $t$
$\tilde{\chi}^{t}=[w_{1}^{t_{1}-\tau_{1}^{t}},\dots,w_{M}^{t_{M}-\tau_{M}^{t}}]$ Delayed learnable parameters of all clients at server time $t$
${\bm{w}}^{t}=[w_{0}^{t},w_{1}^{t_{1}},\dots,w_{M}^{t_{M}}]$ Latest learnable parameters of all clients and the server at server time $t$
$\tilde{{\bm{w}}}^{t}=[w_{0}^{t-\tau_{1}^{t}},w_{1}^{t_{1}-\tau_{1}^{t}},\dots,w_{M}^{t_{M}-\tau_{M}^{t}}]$ Delayed learnable parameters of all clients and the server at server time $t$
$h_{m,i}^{t_{m}\pm}=h_{m}(w_{m}^{t_{m}}\pm\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})$ Local embeddings of client $m$ for data sample $i$ at client time $t_{m}$ under the perturbed parameters
$\delta_{m,i}^{t,t_{m}}$ as defined in (7) zeroth-order difference information from client $m$ and data sample $i$ at server time $t$ and client time $t_{m}$
$g^{t}_{m,i}(\tilde{{\bm{w}}}^{t})=\delta_{m,i}^{t,t_{m}}{\bm{u}}_{m}^{t_{m}}$ zeroth-order gradient estimator from client $m$ and data sample $i$ at server time $t$ (with delay)
$\breve{g}^{t}_{m,i}(\tilde{{\bm{w}}}^{t})=\mathrm{clip}_{C}\left(\delta_{m,i}^{t,t_{m}}\right){\bm{u}}_{m}^{t_{m}}$ Clipped zeroth-order gradient estimator from client $m$ and data sample $i$ at server time $t$ (with delay)
$\breve{G}^{t}_{m}(\tilde{{\bm{w}}}^{t})=(1/B)\sum_{i\in{\mathcal{I}}_{m}^{t_{m}}}\breve{g}^{t}_{m,i}(\tilde{{\bm{w}}}^{t})+z_{m}^{t_{m}}{\bm{u}}_{m}^{t_{m}}$ Clipped differential private zeroth-order gradient estimator from client $m$ at server time $t$ (with delay)
$G^{t}_{m}(\tilde{{\bm{w}}}^{t})=(1/B)\sum_{i\in{\mathcal{I}}_{m}^{t_{m}}}g^{t}_{m,i}(\tilde{{\bm{w}}}^{t})+z_{m}^{t_{m}}{\bm{u}}_{m}^{t_{m}}$ Non-clipped differential private zeroth-order gradient estimator from client $m$ at server time $t$ (with delay)
Table 1: Table of Notations
Note that in the notation table, we use “$\breve{\phantom{x}}$" to define clipped gradient estimators, and we use “$\tilde{\phantom{x}}$" to denote delayed model parameters. In the rest of the proof, we also use gradient estimators parameterized by the no delaying parameters ${\bm{w}}$ instead of $\tilde{{\bm{w}}}$ to assume that we update the model without delay.
To begin with, we restate the assumptions required for establishing the convergence analysis.
Assumption A.1 ($\ell$-Lipschitz).
Assumption A.2 ($L$-Smooth).
[…]
Assumption A.3 (Bounded gradient variance).
Assumption A.4 (Independent Participation).
The probability of one client participating in one communication round is independent of other clients and satisfies
(client $\mathbb{P}(\text{client }m\text{ uploading})=q_{m}$
One of the important parts in the proof of Theorem 5.4 is to bound the zeroth-order gradient estimator. We first introduce the formal definition of zeroth-order two-point gradient estimator that is used in our algorithm, and prove some technical lemmas that reveal some important properties.
Definition A.5.
Lemma A.6.

### v3

@@ V-A Convergence Guarantee and Discussion
Discussion. The first term is influenced by the model’s initialization, $F^{0}$. This term also enjoys the same rate as ZO optimization in the centralized case [ghadimi2013stochastic]. The second term $\mathcal{O}(\frac{d_{*}}{T})$ is a standard term for ZOO methods based on the usage of ZO estimator updates. The third term captures the impact of various noise sources in the learning system. Here, $\sigma_{s}^{2}/B$ represents the noise introduced by stochastic gradients, and $\sigma_{dp}^{2}$ corresponds to the variance of the injected DP noise. While reducing $\sigma_{dp}^{2}$ improves utility, it comes at the cost of weakening privacy guarantees. Consequently, this term encapsulates the fundamental trade-off between model performance, computational cost, and privacy budget. The fourth term, quantifies the impact of the gradient clipping operation on convergence. As $C$ increases, increases, this term diminishes, effectively recovering the non-private case in the limit. However, the noise variance $\sigma_{dp}^{2}$ also scales linearly with $C$ , which can impede overall convergence. This trade-off underscores the importance of carefully tuning the sensitivity level to balance privacy preservation and learning efficiency.
By combining the convergence analysis in Theorem V.4 with the privacy mechanism in Section VI, we obtain the following corollary, which characterizes the privacy-utility tradeoff of DPZV.
Corollary V.5.
Under the assumptions of Theorem V.4, consider the differentially private mechanism described in Section VI, where the variance of the injected DP noise is chosen as $\sigma_{\mathrm{dp}}=\frac{2C\sqrt{T}}{D\mu}.$ If the total number of global iterations is selected as
$T={\mathcal{O}}\!\left(\frac{D^{2}\mu^{2}\sqrt{d}}{C^{2}}\left({\mathcal{F}}+\frac{\sigma^{2}}{B}+\frac{C^{2}}{e^{C^{2}}-1}\right)\right),$ (11)
then the convergence bound in Theorem V.4 implies
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\left\|\nabla\mathcal{F}(w^{t})\right\|^{2}\right]={\mathcal{O}}\!\left(\frac{\sqrt{d}}{D\mu}\right).$ (12)
Discussion. Corollary V.5 shows that, under differential privacy constraints, the proposed zeroth-order method attains the same asymptotic privacy–utility tradeoff as differentially private first-order methods, such as DP-SGD [chen2020understanding]. This result highlights an important contrast with the non-private setting: although zeroth-order optimization generally exhibits slower convergence than first-order methods, the proposed DP noise injection strategy reshapes the dominant error terms in the convergence analysis, allowing zeroth-order methods to achieve comparable asymptotic guarantees.
V-B Proof of Theoretical Results
In this part, we briefly describe the main steps behind the main convergence theorem V.4.
The proof of the theorem follows the standard smoothness- based descent framework for non-convex optimization. However, the convergence is affected by stochastic sampling, zeroth-order perturbations, differential privacy noise, gradient clipping and communication delays, and we need to bound them. The following lemmas solve these questions, and the full details are relegated to Appendix -A.
We first introduce the following lemma, which bound how much the ZO estimator is biased from the ground truth gradient.
Lemma V.6.